The sum of three numbers is 44. The second number is is three times the first number and the third number is 6 less than the first number. Find the three numbers. How do I write this in equations.
Answer by mananth(16946)

let first number be x
second number is 3 times = 3x
third number =x-6
..
add them up
..
x+3x+x-6 = 44
5x-6 =44
+6 +6
5x= 50
/5
x= 50/5
x= 10 first number
3x is the second number = 3*10=30
x-6 is the third number = 3=10-6=4
..
10, 30, 4
..
m.ananth@hotmail.ca+